happypotato:

Find the common ratio and the next term for the sequence shown. Express your answers as fractions or as decimals. 333 1/3, 250, 187 1/2

happypotato:

1/3, 1/2 are fractions

SemiDefinite:

calculate the common ratios

Extrinix:

so, to find the common ratio of \(333\dfrac{1}{3},~250,~187\dfrac{1}{2}\) you divide from the highest number to the number below it, so \(\dfrac{333\dfrac{1}{3}}{250}\) then you find the answer of that, then repeat it for the next set \(\dfrac{250}{187\dfrac{1}{2}}\) now what is the common difference between the numbers? first set is \(1.\overline{3}\) and second set is \(0.\overline{3}\) now, if you noticed, the ratio isn't similar, meaning there is no common ratio. this doesn't mean you can't find the next term though, to find the term you subtract your smallest number from the one above it, so \(250-187\dfrac{1}{2}\) giving you 62.5, now, subtract that from your lowest number \(187\dfrac{1}{2}-62\dfrac{1}{2}\) the halves cancel out, so you're left with \(187-62\) and that is 125, your expected next number.